### v2

# Metaplectic time-frequency representations

Abstract.
Time-frequency representations stemmed in 1932 with the introduction of the Wigner distribution. For most of the 20th century, research in this area primarily focused on defining joint probability distributions for position and momentum in quantum mechanics. Applications to electrical engineering were soon established with the seminal works of Gabor and the researchers at Bell Labs. In 2012, Bai, Li and Cheng used for the first time metaplectic operators, defined in the middle of 20th century by Van Hove, to generalize the Wigner distribution and unify effectively the most used time-frequency representations under a common framework. This work serves as a comprehensive up-to-date survey on time-frequency representations defined by means of metaplectic operators, with particular emphasis on the recent contributions by Cordero and Rodino, who exploited metaplectic operators to their limits to generalize the Wigner distributions. Their idea provides a fruitful framework where properties of time-frequency representations can be explained naturally by the structure of the symplectic group.
Key words and phrases:
Time-frequency analysis, metaplectic representations, linear canonical transforms, Wigner distribution, Modulation spaces
2020 Mathematics Subject Classification:
42C15, 42B35, 42A38
Dedicated to Luigi Rodino as a tribute to his mathematical career.
1. Notation
We denote by $x\xi=x\cdot\xi$ the standard inner product in $\mathbb{R}^{d}$. The notation $\langle\cdot,\cdot\rangle$ is used either to denote the sesquilinear inner product of $L^{2}(\mathbb{R}^{d})$ and its unique extension to a duality pairing (antilinear in the second component) $\mathcal{S}^{\prime}(\mathbb{R}^{d})\times\mathcal{S}(\mathbb{R}^{d})$, where $\mathcal{S}(\mathbb{R}^{d})$ is the Schwartz class of rapidly decreasing smooth functions. If $f,g$ are functions on $\mathbb{R}^{d}$, $f\asymp g$ means that there exist $A,B>0$ such that $Af(x)\leq g(x)\leq Bf(x)$ holds for every $x\in\mathbb{R}^{d}$. If $f,g\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$ are tempered distributions, $\bar{f}$ denotes the complex conjugate of $f$, whereas $f\otimes g(x,y):=f(x)g(y)$ is their tensor product. We work with complex valued functions and distributions, and real matrices. The space of $m\times n$ real matrices is denoted by $\mathbb{R}^{m\times n}$. If $M\in\mathbb{R}^{m\times n}$, $M^{T}\in\mathbb{R}^{n\times m}$ denotes its transpose. If $M$ is a square matrix, $\det(M)$ is its determinant, and $\operatorname*{GL}(d,\mathbb{R})$ is the group of $d\times d$ invertible matrices. We write $M\in\operatorname*{Sym}(d,\mathbb{R})$ if $M$ is a $d\times d$ symmetric matrix, i.e., $M^{T}=M$. $I_{d}$ denotes the $d\times d$ identity matrix, whereas $0_{d}$ denotes the $d\times d$ matrix with all zero entries.
2. Introduction
2.1. The 20th century
Since the seminal work by Wigner [45], published in 1932, where the Wigner distribution
(1) $Wf(x,\xi)=\int_{\mathbb{R}^{d}}f(x+t/2)\overline{f(x-t/2)}e^{-2\pi i\xi t}dt,\qquad f\in L^{2}(\mathbb{R}^{d}),\quad x,\xi\in\mathbb{R}^{d}$
is defined in the context of quantum mechanics, a broad spectrum of time-frequency representations emerged in the literature. According to [45], the Wigner distribution was found by Wigner himself and Szilard some years before 1932 for another, unspecified, purpose. In 1948 Ville introduced the Wigner distribution in the context of signal analysis [43], alongside another time-frequency representation
(2) $Af(x,\xi)=\int_{\mathbb{R}^{d}}f(t+x/2)\overline{f(t-x/2)}e^{-2\pi i\xi t}dt,\qquad f\in L^{2}(\mathbb{R}^{d}),\quad x,\xi\in\mathbb{R}^{d},$
named ambiguity function after the work by Woodward [46], who applied it to radar analysis. Just one year after the publication of Wigner’s work, Kirkwood defined
(3) $W_{0}f(x,\xi)=f(x)\overline{\hat{f}(\xi)}e^{-2\pi i\xi x},\qquad f\in L^{2}(\mathbb{R}^{d}),\quad x,\xi\in\mathbb{R}^{d},$
[31], that was named after Rihacek due to his contribution in 1968 [40]. Parallel to these works, Gabor proposed time-frequency shifts
(4) $\pi(x,\xi)g(t)=e^{2\pi i\xi t}g(t-x),\qquad x,\xi\in\mathbb{R}^{d},$
to decompose signals into fundamental atoms, where he considered a Gaussian window $g$, [25]. The term short-time Fourier analysis emerged naturally over the decades, and the very first appearance of the short-time Fourier transform (STFT)
(5) $V_{g}f(x,\xi)=\langle f,\pi(x,\xi)g\rangle=\int_{\mathbb{R}^{d}}f(t)\overline{g(t-x)}e^{-2\pi i\xi t}dt,\qquad f,g\in L^{2}(\mathbb{R}^{d}),\quad x,\xi\in\mathbb{R}^{d}$
is lost in the history, even though this phraseology was known to Flanagan and Golden, who used it in 1966 in the context of synthesized human speech, [24]. A more experimental evolution occurred for the spectrogram, stemmed from the empirical works of the Bell Labs [32], summarized in the book by Potter, Kopp and Green [38], and later developed in [35, 39]. Specifically, the spectrogram is defined as
(6) ${\operatorname*{Spec}}_{g}(f)=|V_{g}f|^{2},\qquad f,g\in L^{2}(\mathbb{R}^{d}).$
A key effort to identify a common thread unifying time-frequency representations is due to Cohen [7], who observed how most of the above-mentioned time-frequency representations can be written as convolutions between the ambiguity function and a distribution, later named kernel by Claasen and Mecklenbräuker [6]. The family of such time-frequency representations is known as Cohen’s class at the present. It is straightforward to observe that the Cohen’s class can be defined in terms of the Wigner distribution, [6], so that a time-frequency representation $Q$ belongs to the Cohen’s class if
(7) $Qf=Wf\ast\sigma,\qquad f\in\mathcal{S}(\mathbb{R}^{d})$
for some $\sigma\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$. Besides the Wigner distribution, the ambiguity function, and the Rihacek distribution, other elements of the Cohen’s class are discussed in [5, 33, 36, 48].
The survey by Cohen [8] and his subsequent work [9] mark the conclusion of the classical era of time-frequency distributions. For the purpose of this work, it is worth to mention how a crucial part of Cohen’s work was relating properties of representations in the Cohen’s class to the structure of the corresponding kernels.
Other than their different properties, the main aspect distinguishing (1) and (2) from (5) and (6) is the presence of a window function $g$ in the latter. In 1980, Szu and Blodgett [41] defined the cross-Wigner distribution
(8) $W(f,g)(x,\xi)=\int_{\mathbb{R}^{d}}f(x+t/2)\overline{g(x-t/2)}e^{-2\pi i\xi t}dt,\qquad f,g\in L^{2}(\mathbb{R}^{d}),\quad x,\xi\in\mathbb{R}^{d}$
and the cross-ambiguity function
(9) $A(f,g)(x,\xi)=\int_{\mathbb{R}^{d}}f(t+x/2)\overline{g(t-x/2)}e^{-2\pi i\xi t}dt,\qquad f,g\in L^{2}(\mathbb{R}^{d}),\quad x,\xi\in\mathbb{R}^{d}.$
According to the authors, the main interest in introducing the window $g$ was in signal decoding: an unknown signal $f$ could be read correctly by the communicating parties, only with the prior knowledge of the window. In 1985, Janssen [30] considered new representations in the Cohen’s class known nowadays as the $\tau$-Wigner distributions
(10) $W_{\tau}f(x,\xi)=\int_{\mathbb{R}^{d}}f(x+\tau t)\overline{f(x-(1-\tau)t)}e^{-2\pi i\xi t}dt,\qquad f,g\in L^{2}(\mathbb{R}^{d}),\quad x,\xi\in\mathbb{R}^{d},$
where $0\leq\tau\leq 1$. Observe that for $\tau=0$, $W_{\tau}=W_{0}$ is the Rihacek distribution, whereas the classical Wigner distribution corresponds to the choice $\tau=1/2$. The cross-$\tau$-Wigner distributions were considered in [3].
2.2. The metaplectic framework
Metaplectic operators stemmed parallel to time-frequency representations. They were defined in 1951 by Van Hove in his Ph.D. thesis [42], but they comprise operators that were discovered long before, we mention fractional Fourier transforms [10, 34] as non-straightforward examples. In one sentence, the metaplectic group $\operatorname*{Mp}(d,\mathbb{R})$ is the double cover of the symplectic group $\operatorname*{Sp}(d,\mathbb{R})$, see [16, Section 1.1] for the precise definition and the notation. Concretely, it is the group of unitary operators on $L^{2}(\mathbb{R}^{d})$ generated by
(11) $i^{-d/2}\mathcal{F}f(\xi)=i^{-d/2}\int_{\mathbb{R}^{d}}f(x)e^{-2\pi i\xi x}dx,$
(12) $i^{m}\mathfrak{T}_{M}f(x)=i^{m}|\det(M)|^{1/2}f(Mx),\qquad M\in\operatorname*{GL}(d,\mathbb{R}),$
(13) $\mathfrak{p}_{Q}f(x)=e^{i\pi Qx\cdot x}f(x),\qquad Q\in\mathbb{R}^{d\times d},\quad Q^{T}=Q,$
where $f\in\mathcal{S}(\mathbb{R}^{d})$, and $m\in\mathbb{Z}$ is the argument of $\det(M)$ (Maslov index). For simplicity, we omit the phase factors $i^{-d/2}$ and $i^{m}$ in the discussion below. The projection $\pi^{Mp}:\hat{S}\in\operatorname*{Mp}(d,\mathbb{R})\to S\in\operatorname*{Sp}(d,\mathbb{R})$ is a group homomorphism with kernel $\{\pm\operatorname*{id}_{L^{2}}\}$.
The interplay between the main time-frequency representations and fractional Fourier transforms was already examined in 2001 by Pei and Ding [37]. However, the idea of generalizing the Wigner distribution using metaplectic operators is first due to Bai, Li and Cheng, who defined the new Wigner–Ville distribution in [1], by replacing $f$ in (1) with $\hat{S}f$, $\hat{S}\in\operatorname*{Mp}(1,\mathbb{R})$.
In 2015, Z. Zhang and Luo generalized the new Wigner–Ville transform further [47], improving its ability of detecting linear frequency modulated signals under low Signal-to-Noise Ratio. Their construction can be formalized in dimension $d\geq 1$ as follows: fixed $\hat{S},\hat{S}_{1},\hat{S}_{2}\in\operatorname*{Mp}(d,\mathbb{R})$, and $\mathfrak{T}_{M_{1/2}}f(x,t)=f(x+t/2,x-t/2)$,
(14) $W_{\hat{S},\hat{S}_{1},\hat{S}_{2}}f=({\operatorname*{id}}_{L^{2}}\otimes\hat{S})\mathfrak{T}_{M_{1/2}}(\hat{S}_{1}f\otimes\overline{\hat{S}_{2}f}),\qquad f,g\in L^{2}(\mathbb{R}^{d}),$
where $^{\operatorname*{id}}_{L^{2}}\otimes\hat{S}$ is the unique metaplectic operator so that $({\operatorname*{id}}_{L^{2}}\otimes\hat{S})(f\otimes g)=f\otimes\hat{S}g$ for $f,g\in L^{2}(\mathbb{R}^{d})$, see [11, Appendix B].
Parallel to the works by Z. Zhang et al, Bayer, Cordero, Gröchenig and Trapasso defined the (cross-)matrix Wigner distributions in 2019 [2, 19]. They observed that a peculiar detail shared by the distributions in (8), (5), (9) and (10) is that they can be written as
(15) $Q(f,g)=\mathcal{F}_{2}\mathfrak{T}_{M}(f\otimes\bar{g}),\qquad f,g\in L^{2}(\mathbb{R}^{d}),$
for suitable $M\in\operatorname*{GL}(2d,\mathbb{R})$ and
(16) $\mathcal{F}_{2}F(x,\xi)=\int_{\mathbb{R}^{d}}F(x,t)e^{-2\pi i\xi t}dt,\qquad F\in\mathcal{S}({\mathbb{R}^{2d}}),\quad x,\xi\in\mathbb{R}^{d}.$
Up to the factor $|\det(M)|^{-1/2}$, matrix Wigner distributions are defined as in (15). Observe that both $\mathcal{F}_{2}$ and $\mathfrak{T}_{M}$ used in (15) are metaplectic operators acting on $L^{2}({\mathbb{R}^{2d}})$ and, therefore, matrix Wigner distributions are intrinsically constructed through metaplectic operators.
2.3. Metaplectic Wigner distributions
The recent contributions by Cordero and Rodino yielded the definition of the $\mathcal{A}$-Wigner distributions in 2022 [17]. Their core observation was that many of the most important time-frequency representations defined in the last two centuries can be written letting a metaplectic operator on $L^{2}({\mathbb{R}^{2d}})$ act on the tensors
(17) $f\otimes\bar{g}(x,y)=f(x)\overline{g(y)},\qquad f,g\in L^{2}(\mathbb{R}^{d}),\quad x,y\in\mathbb{R}^{d},$
as outlined in the following examples.
Example 2.1.
  1. (i)
    The STFT can be written as
    (18) $V_{g}f=\widehat{\mathcal{A}_{st}}(f\otimes\bar{g}),\qquad f,g\in L^{2}(\mathbb{R}^{d}).$
    The projection of $\widehat{\mathcal{A}_{st}}$ can be computed explicitly as
    (19) $\mathcal{A}_{st}=\begin{pmatrix}I_{d}&-I_{d}&0_{d}&0_{d}\\ 0_{d}&0_{d}&I_{d}&I_{d}\\ 0_{d}&0_{d}&0_{d}&-I_{d}\\ -I_{d}&0_{d}&0_{d}&0_{d}\end{pmatrix}$.
  2. (ii)
    Similarly, $W_{\tau}(f,g)=\widehat{\mathcal{A}_{\tau}}(f\otimes\bar{g})$, where
    (20) $\mathcal{A}_{\tau}=\begin{pmatrix}(1-\tau)I_{d}&\tau I_{d}&0_{d}&0_{d}\\ 0_{d}&0_{d}&\tau I_{d}&-(1-\tau)I_{d}\\ 0_{d}&0_{d}&I_{d}&I_{d}\\ -I_{d}&I_{d}&0_{d}&0_{d}\end{pmatrix}$.
  3. (iii)
[…]
    (21) $S=\begin{pmatrix}A&B\\ C&D\end{pmatrix},\quad\text{and}\quad S_{j}=\begin{pmatrix}A_{j}&B_{j}\\ C_{j}&D_{j}\end{pmatrix}$,
    then $W_{\hat{S},\hat{S}_{1},\hat{S}_{2}}(f,g)=\widehat{\mathcal{A}_{\hat{S},\hat{S}_{1},\hat{S}_{2}}}(f\otimes\bar{g})$, where $W_{\hat{S},\hat{S}_{1},\hat{S}_{2}}(f,g)$ is the cross-version of (14), and
[…]
  4. (iv)
    Other time-frequency representations that were not discussed in the introduction can be interpreted similarly. This is the case of $\hbar$ – STFTs, discussed implicitly in [20]. For $\hbar>0$,
    (23) $V^{\hbar}_{g}f(x,\xi)=(2\pi\hbar)^{-d/2}e^{2\pi i\frac{x\xi}{4\pi\hbar}}V_{g}f(x,\frac{\xi}{2\pi\hbar}),\qquad f,g\in L^{2}(\mathbb{R}^{d}),\quad x,\xi\in\mathbb{R}^{d}$
    (the constant $(2\pi\hbar)^{-d/2}$ is added here to retrieve a unitary rescaling). Then, $V^{\hbar}_{g}f=\widehat{\mathcal{A}_{\hbar}}(f\otimes\bar{g})$ and its projection is the $4d\times 4d$ symplectic matrix
    (24) $\mathcal{A}_{\hbar}=\begin{pmatrix}I_{d}&-I_{d}&0_{d}&0_{d}\\ 0_{d}&0_{d}&2\pi\hbar I_{d}&2\pi\hbar I_{d}\\ 0_{d}&0_{d}&I_{d}/2&-I_{d}/2\\ -\frac{1}{4\pi\hbar}I_{d}&-\frac{1}{4\pi\hbar}I_{d}&0_{d}&0_{d}\end{pmatrix}$.
[…]
Definition 2.2.
For a metaplectic operator $\hat{\mathcal{A}}\in\operatorname*{Mp}(2d,\mathbb{R})$ with projection
(25) $\mathcal{A}=\begin{pmatrix}A_{11}&A_{12}&A_{13}&A_{14}\\ A_{21}&A_{22}&A_{23}&A_{24}\\ A_{31}&A_{32}&A_{33}&A_{34}\\ A_{41}&A_{42}&A_{43}&A_{44}\end{pmatrix}$,
the metaplectic Wigner distribution, or $\mathcal{A}$-Wigner distribution, $W_{\mathcal{A}}$ is the time-frequency representation
(26) $W_{\mathcal{A}}(f,g)=\hat{\mathcal{A}}(f\otimes\bar{g}),\qquad f,g\in L^{2}(\mathbb{R}^{d}).$
Observe that, since $\mathcal{A}$ only determines $\hat{\mathcal{A}}$ up to a phase factor, writing $W_{\mathcal{A}}$ instead of $W_{\hat{\mathcal{A}}}$ is an abuse of notation. As a direct consequence of the continuity properties of metaplectic operators, we have the following result.
Theorem 2.3.
Let $W_{\mathcal{A}}$ be a metaplectic Wigner distribution.
  1. (i)
    $W_{\mathcal{A}}:\mathcal{S}(\mathbb{R}^{d})\times\mathcal{S}(\mathbb{R}^{d})\to\mathcal{S}({\mathbb{R}^{2d}})$ is continuous.
  2. (ii)
    $W_{\mathcal{A}}:L^{2}(\mathbb{R}^{d})\times L^{2}(\mathbb{R}^{d})\to L^{2}({\mathbb{R}^{2d}})$ is bounded and Moyal’s identity holds
    (27) $\langle W_{\mathcal{A}}(f_{1},g_{1}),W_{\mathcal{A}}(f_{2},g_{2})\rangle=\langle f_{1},f_{2}\rangle\overline{\langle g_{1},g_{2}\rangle},\qquad f_{1},f_{2},g_{1},g_{2}\in L^{2}(\mathbb{R}^{d}).$
  3. (iii)
    $W_{\mathcal{A}}:\mathcal{S}^{\prime}(\mathbb{R}^{d})\times\mathcal{S}^{\prime}(\mathbb{R}^{d})\to\mathcal{S}^{\prime}({\mathbb{R}^{2d}})$ is continuous.
3. Main aspects
3.1. Cohen’s class
Examples of metaplectic Wigner distributions belonging to the Cohen’s class are the $\tau$-Wigner distributions. It turns out that the property of belonging to the Cohen’s class is strictly related to covariance, as it is proved in [17, 18].
Definition 3.1.
A metaplectic Wigner distribution $W_{\mathcal{A}}$ is covariant if for every $z\in{\mathbb{R}^{2d}}$,
(28) $W_{\mathcal{A}}(\pi(z)f,\pi(z)g)=W_{\mathcal{A}}(f,g)(\cdot-z),\qquad f,g\in L^{2}(\mathbb{R}^{d}).$
Covariant metaplectic Wigner distributions are characterized in [18, Proposition 2.10] in terms of their projection.
Theorem 3.2.
Let $W_{\mathcal{A}}$ be any metaplectic Wigner distribution having projection $\mathcal{A}\in\operatorname*{Sp}(2d,\mathbb{R})$ with block decomposition (25). The following statements are equivalent.
  1. (i)
    $W_{\mathcal{A}}$ is covariant.
  2. (ii)
    We have that $A_{13},A_{21}\in\operatorname*{Sym}(d,\mathbb{R})$ and
    (29) $\mathcal{A}=\begin{pmatrix}A_{11}&I_{d}-A_{11}&A_{13}&A_{13}\\ A_{21}&-A_{21}&I_{d}-A_{11}^{T}&-A_{11}^{T}\\ 0_{d}&0_{d}&I_{d}&I_{d}\\ -I_{d}&I_{d}&0_{d}&0_{d}\end{pmatrix}$.
  3. (iii)
    $W_{\mathcal{A}}(f,g)=W(f,g)\ast\mathcal{F}^{-1}\Phi_{-B_{\mathcal{A}}}$, where
    (30) $B_{\mathcal{A}}=\begin{pmatrix}A_{13}&I_{d}/2-A_{11}\\ I_{d}/2-A_{11}^{T}&-A_{21}\end{pmatrix}$.
3.2. Shift-invertiblity
Other important characters in time-frequency analysis are modulation spaces, stemmed by the works of Feichtinger [21, 23]. We refer to [16, 28] as exhaustive references for the theory of modulation spaces quickly discussed below. Let $m$ be a $v$-moderate weight on $\mathbb{R}^{2d}$}, $0<p,q\leq\infty$ and $g\in\mathcal{S}(\mathbb{R}^{d})\setminus\{0\}$. We set $\lVert f\rVert_{M^{p,q}_{m}}:=\lVert V_{g}f\rVert_{L^{p,q}_{m}}$, where $L^{p,q}_{m}({\mathbb{R}^{2d}})$ are the mixed-norm Lebesgue spaces, and $V_{g}f$ is the STFT in (5). Different windows $g$ yield to equivalent quasi-norms. In this section, we discuss which metaplectic Wigner distributions $W_{\mathcal{A}}$, other than the STFT, satisfy
(31) $\lVert W_{\mathcal{A}}(f,g)\rVert_{L^{p,q}_{m}}\asymp\lVert f\rVert_{M^{p,q}_{m}}.$
For a symplectic matrix $\mathcal{A}\in\operatorname*{Sp}(2d,\mathbb{R})$ with blocks (25), we identify the submatrix
(32) $E_{\mathcal{A}}=\begin{pmatrix}A_{11}&A_{13}\\ A_{21}&A_{23}\end{pmatrix}$.
Let $f\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$ and $g_{1},g_{2}\in\mathcal{S}(\mathbb{R}^{d})\setminus\{0\}$. The computations in [18, Theorem 2.22] show that for $1\leq p\leq\infty$ and $v_{s}=(1+|\cdot|^{2})^{s/2}$, $s\geq 0$,
(33) $\lVert f\rVert_{M^{p}_{v_{s}}}\lesssim\lVert|W_{\mathcal{A}}(f,g_{1})|\ast|W_{\mathcal{A}}(g_{1},g_{2})|(-E_{\mathcal{A}}\cdot)\rVert_{L^{p}_{v_{s}}}.$
If $E_{\mathcal{A}}$ is invertible, then
(34) $\lVert|W_{\mathcal{A}}(f,g_{1})|\ast|W_{\mathcal{A}}(g_{1},g_{2})|(-E_{\mathcal{A}}\cdot)\rVert_{L^{p}_{v_{s}}}\asymp\lVert|W_{\mathcal{A}}(f,g_{1})|\ast|W_{\mathcal{A}}(g_{1},g_{2})(-\cdot)|\rVert_{L^{p}_{v_{s}}},$
since $v_{s}\asymp v_{s}\circ E_{\mathcal{A}}$. By Young’s inequality,
(35) $\lVert f\rVert_{M^{p}_{v_{s}}}\lesssim\lVert W_{\mathcal{A}}(f,g_{1})\rVert_{L^{p}_{v_{s}}}\lVert W_{\mathcal{A}}(g_{1},g_{2})\rVert_{L^{1}_{v_{s}}},$
and we get the upper bound
(36) $\lVert f\rVert_{M^{p}_{v_{s}}}\lesssim\lVert W_{\mathcal{A}}(f,g_{1})\rVert_{L^{p}_{v_{s}}}.$
In [18] the authors conjectured condition $E_{\mathcal{A}}\in\operatorname*{GL}(d,\mathbb{R})$ to be crucial in the identification of those metaplectic Wigner distributions so that (31) holds.
Definition 3.3.
We say that $W_{\mathcal{A}}$, or equivalently $\mathcal{A}$, is shift-invertible if $E_{\mathcal{A}}\in\operatorname*{GL}(2d,\mathbb{R})$.
Shift-invertibility was discovered for matrix Wigner distributions in [19], but it was formalized only in [18] in the broader context of $\mathcal{A}$-Wigner distributions.
Example 3.4.
  1. (i)
    The Wigner distribution and the STFT are shift-invertible. Moreover, the $\tau$-Wigner distributions are shift-invertible, for every $\tau\neq 0,1$.
  2. (ii)
    Shift-invertible matrix Wigner distributions were studied in [14].
  3. (iii)
    A distribution $W_{\hat{S},\hat{S}_{1},\hat{S}_{2}}$ in (14) is shift-invertible if and only if the projection $S$ of $\hat{S}$ is free. Indeed, a straightforward computation shows that
    (37) $E_{\mathcal{A}_{\hat{S},\hat{S}_{1},\hat{S}_{2}}}=\begin{pmatrix}\frac{1}{2}A_{1}&\frac{1}{2}B_{1}\\ AA_{1}+\frac{1}{2}BC_{1}&AB_{1}+\frac{1}{2}BD_{1}\end{pmatrix}=\frac{1}{2}\begin{pmatrix}I_{d}&0_{d}\\ -2A&B\end{pmatrix}S_{1},$
    which is invertible if and only if $B\in\operatorname*{GL}(d,\mathbb{R})$.
In [11], the authors proved that shift-invertibility was sufficient (minor requirements aside) for (31) to hold for metaplectic Wigner distributions in the Banach setting, by direct estimating $\lVert W_{\mathcal{A}}(f,g)\rVert_{L^{p,q}_{m}}$ from above and below. Later, in [13], the same authors retrieved the characterization of shift-invertible distributions as rescaled STFTs. To be exhaustive, if $\mathcal{A}$ has blocks (25), we need to define the symmetric matrix
(38) $M_{\mathcal{A}}=\begin{pmatrix}A_{11}^{T}A_{31}+A_{21}^{T}A_{41}&A_{11}^{T}A_{33}+A_{21}^{T}A_{43}\\ A_{13}^{T}A_{31}+A_{23}^{T}A_{41}+I_{d}&A_{13}^{T}A_{33}+A_{23}^{T}A_{43}\end{pmatrix}$,
and the symplectic matrix (see [13])
(39) $G_{\mathcal{A}}=\begin{pmatrix}0_{d}&I_{d}\\ I_{d}&0_{d}\end{pmatrix}E_{\mathcal{A}}^{-1}\begin{pmatrix}A_{12}&A_{14}\\ A_{22}&A_{24}\end{pmatrix}$.
Proposition 3.5.
Let $W_{\mathcal{A}}$ be a metaplectic Wigner distribution, and $\mathcal{A}$ be the corresponding projection having blocks (25). The following statements are equivalent.
  1. (i)
    $W_{\mathcal{A}}$ is shift-invertible.
  2. (ii)
    For every $f,g\in L^{2}(\mathbb{R}^{d})$,
    (40) $W_{\mathcal{A}}(f,g)(z)=c_{\mathcal{A}}|\det(E_{\mathcal{A}})|^{-1/2}\Phi_{M_{\mathcal{A}}}(E_{\mathcal{A}}^{-1}z)V_{\widehat{\delta_{\mathcal{A}}}g}f(E_{\mathcal{A}}^{-1}z),\qquad z\in{\mathbb{R}^{2d}},$
    for some constant $c_{\mathcal{A}}\in\mathbb{C}$, $|c_{\mathcal{A}}|=1$, where $\widehat{\delta_{\mathcal{A}}}=\mathcal{F}\widehat{\overline{G_{\mathcal{A}}}}$.
  3. (iii)
    There exist $E\in\operatorname*{GL}(2d,\mathbb{R})$, $C\in\operatorname*{Sym}(2d,\mathbb{R})$, $\hat{\delta}\in\operatorname*{Mp}(d,\mathbb{R})$ and $c\in\mathbb{C}$, $|c|=1$, such that
    (41) $W_{\mathcal{A}}(f,g)(z)=c|\det(E)|^{1/2}\Phi_{C}(Ez)V_{\widehat{\delta}g}f(Ez),\qquad f,g\in L^{2}(\mathbb{R}^{d}),\quad z\in{\mathbb{R}^{2d}}.$
The set of $4d\times 4d$ shift-invertible symplectic matrices is studied in [12] and [26].
Theorem 3.6.
Let $W_{\mathcal{A}}$ be a metaplectic Wigner distribution, $\mathcal{A}$ be the corresponding projection having blocks (25), and $E_{\mathcal{A}}$ be the submatrix of $\mathcal{A}$ defined as in (32). Let the $v$-moderate weight $m$ satisfy $m\circ E_{\mathcal{A}}\asymp m$.
  1. (i)
    The equivalence (31) holds for every $0<p=q\leq\infty$.
  2. (ii)
    If $E_{\mathcal{A}}$ is upper block triangular ($A_{21}=0_{d}$), then (31) holds for every $0<p,q\leq\infty$.
Moreover, if $p,q\geq 1$, the window $g$ in (31) can be chosen in $M^{1}_{v}(\mathbb{R}^{d})\setminus\{0\}$.
A result similar to Theorem 3.6 can be formulated for Wiener amalgam spaces $W(\mathcal{F}L^{p}_{m_{1}},L^{q}_{m_{2}})(\mathbb{R}^{d})=\mathcal{F}(M^{p,q}_{m_{1}\otimes m_{2}}(\mathbb{R}^{d}))$, see [11, Corollary 3.12] and [13, Corollary 7.2]. These spaces were defined by Feichtinger in 1981 [22].
Theorem 3.6 tells that shift-invertibility is sufficient, along with other minor conditions, to characterize modulation spaces. On the contrary, the Rihacek distribution is not shift-invertible with
(42) $\lVert W_{0}(f,g)\rVert_{L^{p,q}}=\lVert f\rVert_{p}\lVert\hat{g}\rVert_{q},$
and right-hand side is not proportional to $\lVert f\rVert_{M^{p,q}}$, unless $p=q=2$. This fact is more general. In [26] it is proved that $M^{p}$ quasi-norms cannot be measured by means of non shift-invertible distributions.
Theorem 3.7.
Let $W_{\mathcal{A}}$ be a non shift-invertible metaplectic Wigner distributions, and $g\in\mathcal{S}(\mathbb{R}^{d})\setminus\{0\}$. Then, if $0<p\leq\infty$, $p\neq 2$,
(43) $\{f\in\mathcal{S}^{\prime}(\mathbb{R}^{d}):\lVert W_{\mathcal{A}}(f,g)\rVert_{{p}}<\infty\}\neq M^{p}(\mathbb{R}^{d}).$
3.3. Generalized spectrograms
Generalized spectrograms were defined by Boggiatto, De Donno and Oliaro in 2007 [4]: for fixed $\varphi,\psi\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$, and $f,g\in\mathcal{S}^{\prime}(\mathbb{R}^{d})$,
(44) ${\operatorname*{Spec}}^{\varphi,\psi}(f,g)=V_{\varphi}f\overline{V_{\psi}g}.$
These distributions are metaplectic Wigner distributions under appropriate choices of $\varphi$ and $\psi$. The reader may find a complete characterization of metaplectic Wigner distributions that define generalized spectrograms in [15, Theorem 4.6].
3.4. Uncertainty principles
Uncertainty principles for metaplectic Wigner distributions have been studied independently by many authors. Most of the known results are restricted to particular instances of these time-frequency representations, where the computations are explicit, see the early survey [27] and [44]. In [29], the authors formulate Benedick’s uncertainty principle for metaplectic Wigner distributions, characterizing which time-frequency representations satisfy
$W_{\mathcal{A}}(f,g)$ supported on a set of finite measure $\Rightarrow$ $f=0$ or $g=0$
in terms of the Iwasawa decomposition of the corresponding projection $\mathcal{A}$.
4. Inversion formula and Gabor frames
The content of this section is contained in [13]. We refer the reader to [16, 28] for the theory of (Gabor) frames. For every metaplectic Wigner distribution $W_{\mathcal{A}}$ there exists a family of continuous operators $\pi_{\mathcal{A}}(z):\mathcal{S}(\mathbb{R}^{d})\to\mathcal{S}^{\prime}(\mathbb{R}^{d})$, $z\in{\mathbb{R}^{2d}}$, so that
(45) $W_{\mathcal{A}}(f,g)(z)=\langle f,\pi_{\mathcal{A}}(z)g\rangle,\qquad f,g\in\mathcal{S}(\mathbb{R}^{d})$
(metaplectic atoms).
Theorem 4.1.
If $f\in L^{2}(\mathbb{R}^{d})$ and $g,\gamma\in\mathcal{S}(\mathbb{R}^{d})$ are so that $\langle g,\gamma\rangle\neq 0$, then
(46) $f\overset{L^{2}}{=}\frac{1}{\langle\gamma,g\rangle}\int_{{\mathbb{R}^{2d}}}W_{g}f(z)\pi_{\mathcal{A}}(z)\gamma dz.$
In general, metaplectic atoms map $\mathcal{S}(\mathbb{R}^{d})$ to $\mathcal{S}^{\prime}(\mathbb{R}^{d})$. However, metaplectic atoms of shift-invertible metaplectic Wigner distributions define surjective quasi-isometries of $L^{2}(\mathbb{R}^{d})$. In this case, we can consider a system $\mathcal{G}_{\mathcal{A}}(g,\Lambda)=\{\pi_{\mathcal{A}}(\lambda)g\}_{\lambda\in\Lambda}$, where $g\in L^{2}(\mathbb{R}^{d})\setminus\{0\}$ and $\Lambda\subset{\mathbb{R}^{2d}}$ is discrete. We say that $\mathcal{G}_{\mathcal{A}}$ is a metaplectic Gabor frame (for $L^{2}(\mathbb{R}^{d})$) if it is a frame for $L^{2}(\mathbb{R}^{d})$. In view of (40), it is straightforward to relate metaplectic Gabor frames to classical Gabor frames, see [13, Theorem 6.4]. In the case of shift-invertible distributions, the deformation operator $\widehat{\delta_{\mathcal{A}}}$ in (40) plays a central role in the frame theory associated to $W_{\mathcal{A}}$. The frame operator associated to $\mathcal{G}_{\mathcal{A}}(g,\Lambda)$ is $S_{\mathcal{A}}f=\sum_{\lambda\in\Lambda}W_{\mathcal{A}}(f,g)(\lambda)\pi_{\mathcal{A}}(\lambda)g$, $f\in L^{2}(\mathbb{R}^{d})$, and the canonical dual window of $g$ is $\gamma_{\mathcal{A}}=\widehat{\delta_{\mathcal{A}}}^{-1}S_{\mathcal{A}}^{-1}\widehat{\delta_{\mathcal{A}}}g$. We have that if $m\asymp m\circ E_{\mathcal{A}}$, then for every $0<p,q\leq\infty$,
(47) $f\overset{M^{p,q}_{m}}{=}\sum_{\lambda\in\Lambda}W_{\mathcal{A}}(f,g)(\lambda)\pi_{\mathcal{A}}(\lambda)\gamma_{\mathcal{A}},$
with unconditional convergence if $\max\{p,q\}\neq\infty$, and weak-$\ast$ convergence in $M^{\infty}_{1/v}$, otherwise. Moreover,
(48) $\lVert f\rVert_{M^{p,q}_{m}}\asymp\lVert(W_{\mathcal{A}}(f,g)(\lambda))_{\lambda\in\Lambda}\rVert_{\ell^{p,q}_{m}},$
where $\ell^{p,q}_{m}(\Lambda)$ are the mixed-norm sequence spaces. We refer the reader to [13, Theorem 7.3] for the precise statement.